# TokenVerse: Unifying Speech and NLP Tasks via Transducer-based ASR

### v1

@@ Abstract
In traditional conversational intelligence from speech, a cascaded pipeline is used, involving tasks such as voice activity detection, diarization, transcription, and subsequent processing with different NLP models for tasks like semantic endpointing and named entity recognition (NER). Our paper introduces TokenVerse, a single Transducer-based model designed to handle multiple tasks. This is achieved by integrating task-specific tokens into the reference text during ASR model training, streamlining the inference and eliminating the need for separate NLP models. In addition to ASR, we conduct experiments on 3 different tasks: speaker change detection, endpointing, and NER. Our experiments on a public and a private dataset show that the proposed method improves ASR by up to 7.7% in relative WER while outperforming the cascaded pipeline approach in individual task performance. Additionally, we present task transfer learning to a new task within an existing TokenVerse.
keywords:
multitask training, speech recognition, speaker change detection, named entity recognition, XLSR-Transducer1 Introduction
@@
Automated analysis of conversational audios has a wide range of practical applications, including in contact center analytics [1, 2]. Traditionally, conversational audios are transcribed with intermediate voice activity detection (VAD) [3] or endpointing [4] and diarization [5]. Afterward, separate NLP pipelines are employed on the transcripts to perform tasks such as named entity recognition (NER) [6], among others, to comprehend the conversation’s structure and content [7, 8]. Using separate models for each subtask (optimized independently) has drawbacks [9] such as error propagation and a potential mismatch between automatic speech recognition (ASR) metrics and the final task. For instance, the best ASR hypothesis may not be optimal for the final task. Moreover, the cascaded approaches could translate to increased compute and latency, which will be exacerbated by the introduction of a new task.
@@
In this paper, we introduce TokenVerse, a neural Transducer [10] model capable of learning ASR and multiple additional tasks through the incorporation of task tokens. In contrast to the multi-head based multitasking approaches explored in previous studies [11, 12, 13], TokenVerse distinguishes itself by generating tokens directly within the ASR hypothesis, as illustrated in Fig. 1a. Leveraging the transducer architecture [10], we can attain text-audio alignment for each output token, including those designated as task tokens. For example, we can perform NER directly in the acoustic domain, presenting potential utility in scenarios such as audio de-identification [14]. To address challenges in low-resource settings, we use self-supervised (SSL) trained XLSR-53 [15] model as an encoder in the transducer setup, leading to the XLSR-Transducer (Fig. 1b). Previous works aims at modeling several tasks directly from speech using special tokens [16, 17], or ASR with speaker change detection (SCD) [18, 19, 13], VAD [20], speech-to-text translation [21], or timestamps [22], NER [9, 23] and multi-speaker ASR [24, 25]. Token-based multitasking offers multiple benefits, e.g., it has a fix number of parameters while all tasks are predicted with standard decoding without increased latency. However, NLP tasks like NER in conjunction with other tasks from audio domains have not received much attention in the literature. Therefore, we consider 3 additional tasks alongside ASR: SCD, endpointing and NER. These tasks are selected to represent both audio and NLP domains. SCD is an audio task [26]. Endpointing can be viewed as an NLP task when conducting semantic endpointing [27], or as an audio task [4]. NER is an NLP task [6, 9]. They serve as suitable benchmarks for evaluating our proposed method.
@@ 2 TokenVerse
Through TokenVerse, we aim to train a single model for ASR (main task), speaker change detection (SCD), endpointing, and named entity recognition (NER). This is achieved by augmenting the reference text, with task tokens that denote special events at the acoustic level. In the following sections, we discuss the annotation protocol, dataset preparation, details of our ASR model and ablation experiments.
@@ 2.1 Token Augmentation Protocol
We introduce ”tokens” for tasks apart from ASR: [SCD] (speaker change detection), [NE] and [/NE] (named entity recognition), and [ENDP] (endpointing). An illustrative example is depicted in Figure 1a. We insert [SCD] token during text concatenation if there is a speaker change from one segment to another within an utterance. The [ENDP] token is inserted at the end of a segment text, considered as a semantic endpoint from the conversational context perspective. Note that occurrence of [ENDP] will be a superset of [SCD] because a speaker change indicates the completion of the previous speaker’s sentence. For NER, we insert [NE] before the start of a named entity (NE) and [/NE] after it is concluded, since it can comprise multiple words.
2.2 Dataset Preparation
Our work is focused on conversational audios which is typically long in duration (avg 5 minutes) and can’t be directly used for ASR training due to high GPU memory requirements. The dataset provides audio-text transcripts together with timestamp information for every segment within the long-form audio. For each sample, we begin with the first segment and find the farthest segment such that the duration is up to 20 seconds. Audios within this range are extracted as one utterance and this procedure is repeated until the last segment is consumed. Note that an utterance may span over multiple segments, potentially containing silences, noise, speaker changes, endpoints and numerous named entities. Afterward, we concatenate the text corresponding to all segments within an utterance, inserting token at appropriate positions according to our tasks, described in §2.1. This multitask dataset preparation approach applies universally across all datasets (see §4.1) used in our experiments.
@@ 2.3 Training & Inference
TokenVerse Training We train the XLSR-Transducer model on the multitask data which consists of XLSR encoder, state-less predictor [28] and joint networks (linear layer). The model is trained with pruned transducer loss [29]. We utilize SentencePiece [30] tokenizer to train subwords from the training text [31]. It is important to note that the text includes task-specific tokens, and splitting them into multiple subwords may degrade their prediction accuracy because the entire sequence of subwords for a token must be predicted correctly to count it as a valid token prediction. Hence, we ensure that tokens are represented by a single subword during their training.111https://github.com/google/sentencepiece
TokenVerse Inference We generate hypothesis with beam search. From the hypothesis, we can extract and align the predicted task tokens in the time domain. Since NER consists of two tokens, we extract words between a matched pair of [NE] and [/NE]. We discard any unpaired tag from the hypothesis. To obtain timestamps for [SCD] or [ENDP], we note the acoustic frame index for which these tokens are emitted and calculate time information, i.e., XLSR acoustic embeddings have a frame duration of 25ms and a stride of 20ms. Particularly for [SCD], the time-level token prediction enables subsequent tasks, e.g., diarization [19].
@@ 2.4 Ablations within TokenVerse
Single task For each task, we retain only the tokens specific to that task in the multitask dataset and train our XLSR-Transducer model. This helps eliminate any detractor tasks that may affect the performance of the task being evaluated and serves as a baseline in this paper.
Leave-one-task-out We systematically exclude tokens corresponding to a single task from the multitask dataset and proceed to train our ASR model These experiments aims to examine how the removal of a task affects all other tasks, including ASR. This provides insights into whether we should retain or discard any task in TokenVerse for optimal performance on a given task.
Task-Transfer Learning In conventional multi-head multitask architectures [11, 12], integrating a new task typically necessitates fine-tuning the model on the specific task while keeping the base encoder and other heads frozen. We explore the viability of this extension for TokenVerse by fine-tuning the model, derived from the removal of a task, specifically on the new task. Furthermore, we evaluate its impact on both existing tasks and the performance of the new task in comparison to the overall performance when all-tasks are included.
@@
3 Task-Specific Baselines, Metrics & Evaluation Protocol
Automatic Speech Recognition We train our XLSR-Transducer model after removing all task tokens from the multitask dataset. This serves as a baseline for comparison with the multitask models on the ASR task. Evaluation It is evaluated with WER. For TokenVerse models, we remove task tokens from both the reference and hypothesis to compute WER for a fair comparison. We also report WER including task tokens, which reflects its prediction errors.
Named-Entity Recognition We finetune pretrained BERT222https://huggingface.co/google-bert/bert-base-uncased [32] model on our datasets for subword-level NER classification. We evaluate the models on both reference and hypothesis from the ASR model. Evaluation NER systems are usually evaluated by comparing their outputs against human annotations, either using an exact-match or soft-match approach [6]. We adapted these metrics to a scenario where the text comes from an ASR system. Exact-Match: Let be the set of predicted entities, and be the set of actual entities, where each and is accompanied by its corresponding [NE]-[/NE] tokens (See Fig.1). Thus, an entity is considered correctly identified if and only if: , including the tokens. Soft-Match: in this case we only count for the paired sets of [NE]-[/NE] tokens without considering if the predicted entity value was correctly transcribed. After obtaining each pair, we evaluate NER with F1-score.
Speaker Change Detection For the SCD baseline, we utilize the diarization pipeline333huggingface.co/pyannote/speaker-diarization-3.1 from PyAnnote [33] to extract speaker change timestamps from the audio. In literature, the SCD is predominantly regarded as a task within the audio domain [26], we opt not to establish an independent text-based baseline for this task. Evaluation We evaluate SCD in two ways: text-based (only valid for TokenVerse) and time-based. In text-based evaluation, we align the reference and hypothesis using edit-distance. For each occurrence of the [SCD] token in the reference, matching with the same token in the hypothesis counts as True Positive; else, False Negative. Unmatched tokens in the hypothesis are considered False Positive. F1 score is calculated by standard definitions. In time-based evaluation, we obtain the timestamps where [SCD] tokens are predicted in the hypothesis. We calculate F1 score [13], using a collar of 250ms during timestamp matching, following common practice in speaker diarization literature [5]. Additionally, segment coverage, purity [26], and their F1 score are also reported. We use pyannote.metrics [34] to compute all time-based metrics.
Endpointing Considering semantic endpointing, we fine-tune BERT [32] for [ENDP] token classification on the multitask training text, termed as BERT-ENDP. Results are reported on both reference text and hypothesis text obtained from TokenVerse. From the audio perspective, we use segmentation pipeline444huggingface.co/pyannote/segmentation-3.0 from PyAnnote to obtain endpoint timestamps. Evaluation Endpointing is also evaluated in two ways: text-based and time-based. The text-based evaluation follows the same approach as described previously for SCD. In the time-based evaluation, the F1 score computation also follows the same approach as for SCD. Additionally, we also report false alarms (FA), missed speech (MS), and detection error rate (DER), which are common metrics in endpointing literature [3].
4 Experimental Setup
4.1 Dataset
@@
To train TokenVerse, we require conversational audio data with corresponding transcripts, NER and segment timestamps, and speaker annotations. We could not find a large-scale public dataset satisfying all the tasks. Thus, we opt for a private dataset (DefinedAI555https://www.defined.ai/) which contains stereo-audio/transcript pairs for contact center conversations between agents and customers. We upsampled audio from 8 kHz to 16 kHz to align with the XLSR-53 model’s requirements. Each segment includes transcripts, speaker ID and NE annotations, facilitating multitask dataset preparation (Sec 2.2). This dataset spans health, banking and finance domains, which makes it particularly challenging due to variations in NEs. Additionally, we train and evaluate TokenVerse on the open-source CallHome English dataset (LDC97S42), which contains natural conversational stereo-audios between multiple speakers. The transcript includes named entities annotation.This dataset poses challenges due to its natural conversational nature, known to be challenging for ASR modeling, and a large number of short segments without entities, differing from the DefinedAI dataset. Further details about these datasets are provided in Table 1.
4.2 Training TokenVerse
@@
We train TokenVerse on the multitask dataset. It involves XLSR-transducer model, which is constructed from the Icefall’s Transducer recipe666https://github.com/k2-fsa/icefall/tree/master/egs/librispeech/ASR/zipformer adapted with XLSR from fairseq [35] as the encoder. The fine-tuning uses Scaled Adam [36] and a learning rate scheduler that consists of a 500-step warmup phase followed by a decay phase directed by the number of steps and epochs. The model is optimized with pruned RNN-t loss [29]. The learning rate is set to and we train the model for 50 epochs. For each dataset, the best epoch is selected based on the WER on respective dev sets and results are presented on the eval sets. The task-transfer experiments, described in §2.4, are trained for additional 10 epochs on the new task.
| Exp | Model | w/ token | w/o token† |
| 1) | ASR (baseline) | 15.3 | |
| 2) | all-tasks | 15.6 | 14.7 |
| 3-a) | single-[SCD] | 15.2 | 15.1 |
| 3-b) | single-[NE] | 15.3 | 14.7 |
| 3-c) | single-[ENDP] | 14.8 | 14.7 |
@@ 5 Results & Discussion
Automatic Speech Recognition For the DefinedAI (Tab. 2) set, WERs are reported both with and without task tokens in the reference and hypothesis for multitask models. However, the baseline ASR model is trained without task tokens in transcripts, so there is no distinction between them. Including all tasks in TokenVerse (exp 2) leads to a 4% relative improvement in WER compared to the baseline ASR model (exp 1). For models trained on a single task (exp 3a-c), ASR results remain similar except for SCD. When comparing WERs before and after token removal, we observe a relatively large gap between all-tasks and single-task models, potentially due to higher token insertion or deletion as compared to non-token words in the hypothesis. In single-task models, a larger gap is observed for [NE] as the model must accurately predict both tokens, introducing additional error sources. On the CallHome dataset (Tab. 5), the multitask model with all tokens yields a 7.7% relative improvement. Overall, the results on both datasets indicate that the all-tasks TokenVerse improves ASR performance.
Named-Entity Recognition As expected, compared to evaluating BERT-NER on reference text, a significant degradation is observed when evaluated on hypothesis (Tab. 3) due to ASR errors [9]. In exact-match, on both the DefinedAI (Tab. 3) and CallHome (Tab. 5) test sets, the all-tasks TokenVerse outperforms the baseline BERT-NER models trained on their respective datasets and evaluated on hypothesis in F1 score. This is not the case for soft-match evaluation on the DefinedAI test set, where the F1 score is similar. This degradation is mostly attributed to the incorrect prediction of [/NE] tag by the baseline, resulting in only a partial match of the named entity words. The absolute F1 score is low on the CallHome dataset due to higher ASR errors on named entities, attributed to their low repetition in the training text (see Tab. 1).
Speaker Change Detection On the DefinedAI (Tab. 4), including all tasks in TokenVerse outperforms the baseline PyAnnote model in time-based evaluations. Interestingly, models trained for single-task SCD perform better than the all-tasks model in terms of F1, but show similar results for Coverage-Purity based F1. Upon closer scrutiny, we found that including [ENDP] delays the prediction for [SCD] tokens, causing the hypothesis timestamps of these tokens to fall outside the tolerance window (250ms). Increasing the tolerance window further improves the F1 for both models, with a much higher rate of increase for the all-tasks model. This observation is reinforced in the text-based F1 score, where the all-tasks model achieves an F1 score of 90.3% compared to 88.5% from the single-[SCD] model. On the CallHome test set (Tab. 5), the all-tasks model outperforms the PyAnnote baseline. These evaluations suggest that excluding [SCD] from TokenVerse is preferable for precise speaker change timestamps, while including all tasks improves speaker-attributed text segmentation.
@@
5.1 Ablation results
In ASR, we observed degradation for all ablation experiments (see §2.4), with the largest relative degradation of 2.4% in WER when [ENDP] was removed. Transfer learning on any of the 3 tasks do not degrade ASR performance further. The text-based evaluations of other tasks on DefinedAI are reported in Figure 2; absolute change is calculated from the all-tasks model. Removing a task adversely affects other tasks. Specifically, for SCD and endpointing, [NE] removal has the least impact on performance. Learning it afterward either improves or maintain their performance, indicating a stronger correlation between these tasks than with NER; supported by the degradation in [SCD] performance when [ENDP] is removed. Task transfer on [ENDP] degrades the performance further, possibly due to confusion during prediction caused by the insertion of the token before [SCD] during training. Transfer to NER shows relatively large degradation compared to other tasks, likely because the model must predict both [NE] and [/NE] accurately. This suggests that tasks encoded with multiple tokens may not transfer as effectively as those encoded with a single token.
Overall, all-tasks TokenVerse outperforms specialized models for each task and single-task models suggesting that additional tasks improve each other. Moreover, our task transfer experiments suggest that a new task can be learned effectively.

### v2

# TokenVerse: Towards Unifying Speech and NLP Tasks via Transducer-based ASR

@@ Abstract
In traditional conversational intelligence from speech, a cascaded pipeline is used, involving tasks such as voice activity detection, diarization, transcription, and subsequent processing with different NLP models for tasks like semantic endpointing and named entity recognition (NER). Our paper introduces TokenVerse, a single Transducer-based model designed to handle multiple tasks. This is achieved by integrating task-specific tokens into the reference text during ASR model training, streamlining the inference and eliminating the need for separate NLP models. In addition to ASR, we conduct experiments on 3 different tasks: speaker change detection, endpointing, and NER. Our experiments on a public and a private dataset show that the proposed method improves ASR by up to 7.7% in relative WER while outperforming the cascaded pipeline approach in individual task performance. Our code is publicly available: https://github.com/idiap/tokenverse-unifying-speech-nlp
TokenVerse: Towards Unifying Speech and NLP Tasks via Transducer-based ASR
Shashi Kumar1,2 Srikanth Madikeri1,3 Juan Zuluaga-Gomez1 Iuliia Thorbecke1,3 Esaú Villatoro-Tello1 Sergio Burdisso1 Petr Motlicek1,4 Karthik Pandia5 Aravind Ganapathiraju5 1Idiap Research Institute, Switzerland; 2EPFL, Switzerland; 3University of Zurich, Switzerland; 4Brno University of Technology, Czech Republic; 5Uniphore, India shashi.kumar@idiap.ch
1 Introduction
@@
Automated analysis of conversational audios has a wide range of practical applications, including in contact center analytics (Saberi et al., 2017; Mamou et al., 2006). Traditionally, conversational audios are transcribed with intermediate voice activity detection (VAD) (Medennikov et al., 2020) or endpointing (Chang et al., 2019) and diarization (Park et al., 2022). Afterward, separate NLP pipelines are employed on the transcripts to perform tasks such as named entity recognition (NER) (Li et al., 2020), among others, to comprehend the conversation’s structure and content (Zou et al., 2021; Xu et al., 2021). Using separate models for each subtask (optimized independently) has drawbacks (Ghannay et al., 2018) such as error propagation and a potential mismatch between automatic speech recognition (ASR) metrics and the final task. For instance, the best ASR hypothesis may not be optimal for the final task. Moreover, the cascaded approaches could translate to increased compute and latency, which will be exacerbated by the introduction of a new task.
@@
In this paper, we introduce TokenVerse, a neural Transducer (Graves, 2012) model capable of learning ASR and multiple additional tasks through the incorporation of task tokens. In contrast to the multi-head based multitasking approaches explored in previous studies (Chen et al., 2021; wen Yang et al., 2021; Kumar et al., 2024), TokenVerse distinguishes itself by generating tokens directly within the ASR hypothesis, as illustrated in Fig. 1a. Leveraging the transducer architecture (Graves, 2012), we can attain text-audio alignment for each output token, including those designated as task tokens. For example, we can perform NER directly in the acoustic domain, presenting potential utility in scenarios such as audio de-identification (Cohn et al., 2019). To address challenges in low-resource settings, we use self-supervised (SSL) trained XLSR-53 (Conneau et al., 2020) model as an encoder in the transducer setup, leading to the XLSR-Transducer (Fig. 1b). Previous works aims at modeling several tasks directly from speech using special tokens (Wu et al., 2024; Chang et al., 2023), or ASR with speaker change detection (SCD) (Shafey et al., 2019; Xia et al., 2022; Kumar et al., 2024), VAD (Radford et al., 2023), speech-to-text translation (Zuluaga-Gomez et al., 2023), or timestamps (Cornell et al., 2023), NER (Ghannay et al., 2018; Yadav et al., 2020) and multi-speaker ASR (Kanda et al., 2022; Wu et al., 2023). Token-based multitasking offers multiple benefits, e.g., it has a fix number of parameters while all tasks are predicted with standard decoding without increased latency. However, NLP tasks like NER in conjunction with other tasks from audio domains have not received much attention in the literature. Therefore, we consider 3 additional tasks alongside ASR: SCD, endpointing and NER. These tasks are selected to represent both audio and NLP domains. SCD is an audio task (Bredin et al., 2017). Endpointing can be viewed as an NLP task when conducting semantic endpointing (Raux and Eskenazi, 2008), or as an audio task (Chang et al., 2019). NER is an NLP task (Li et al., 2020; Ghannay et al., 2018). They serve as suitable benchmarks for evaluating our proposed method.
@@ 2 TokenVerse
Through TokenVerse, we aim to train a single model for ASR (main task), speaker change detection (SCD), endpointing, and named entity recognition (NER). This is achieved by augmenting the reference text, with task tokens that denote special events at the acoustic level.
@@ 2.1 Token Augmentation Protocol
We introduce “tokens" for tasks apart from ASR: [SCD] (speaker change detection), [NE] and [/NE] (named entity recognition), and [ENDP] (endpointing) to prepare the multitask dataset. An illustrative example is depicted in Figure 1a. We insert [SCD] token during text concatenation if there is a speaker change within an utterance. The [ENDP] token is inserted at the end of a segment text, considered as a semantic endpoint from the conversational context perspective. Note that occurrence of [ENDP] will be a superset of [SCD] because a speaker change indicates the completion of the previous speaker’s sentence. For NER, we insert [NE] before the start of a named entity and [/NE] after it is concluded, since it can comprise multiple words.
@@ 2.2 Training & Inference
TokenVerse Training We train the XLSR-Transducer model on the multitask data which consists of XLSR encoder, state-less predictor (Ghodsi et al., 2020) and joint networks (linear layer). The model is trained with pruned transducer loss (Kuang et al., 2022). We utilize SentencePiece (Kudo and Richardson, 2018) tokenizer to train subwords from the training text (Sennrich et al., 2016). Note that the text includes task-specific tokens, and splitting them into multiple subwords may degrade their prediction accuracy because the entire sequence of subwords for a token must be predicted correctly to count it as a valid token prediction. Hence, we ensure that tokens are represented by a single subword during their training.111https://github.com/google/sentencepiece
TokenVerse Inference We generate hypothesis with beam search. From the hypothesis, we can extract and align the predicted task tokens in the time domain. Since NER consists of two tokens, we extract words between a matched pairs of [NE] and [/NE]. To obtain timestamps for [SCD] or [ENDP], we note the acoustic frame index for which these tokens are emitted and calculate time information, i.e., XLSR acoustic embeddings have a frame duration of 25ms and a stride of 20ms. Particularly for [SCD], the time-level token prediction enables subsequent tasks, e.g., diarization (Xia et al., 2022).
@@ 2.3 Ablations within TokenVerse
Single task For each task, we retain only the tokens specific to that task in the multitask dataset and train our ASR model. This eliminates any detractor tasks that may affect the task being evaluated and serves as a baseline in this paper.
Leave-one-task-out We exclude tokens of a single task from the multitask data and train our ASR model. This provides insights whether we should retain or discard any task in TokenVerse for optimal performance on a given task.
Task-Transfer Learning In multi-head multitask architectures (Chen et al., 2021), a new task can be learnt by fine-tuning the model on the new task while keeping the base encoder and other heads frozen. We explore this for TokenVerse by fine-tuning the model, derived from the removal of a task, on the removed task. Furthermore, we evaluate its impact on both existing tasks and the performance of the new task in comparison to the overall performance when all-tasks are included.
3 Task-Specific Baselines, Metrics & Evaluation Protocol
Automatic Speech Recognition We train our XLSR-Transducer model after removing all task tokens from the multitask dataset. This serves as a baseline for the ASR task. Evaluation It is evaluated with WER. For TokenVerse models, we remove task tokens from both the reference and hypothesis to compute WER for a fair comparison.
Named-Entity Recognition We finetune pretrained BERT222https://huggingface.co/google-bert/bert-base-uncased (Devlin et al., 2019) model on our datasets for subword-level NER classification, a commonly used approach for this task in the literature. We evaluate the models on both reference and hypothesis from the ASR model. Evaluation NER systems are usually evaluated by comparing their outputs against human annotations, either using an exact-match or soft-match approach (Li et al., 2020). We adapted these metrics to a scenario where the text comes from an ASR system. Detailed description in appendix A.
Speaker Change Detection We utilize the diarization pipeline333https://huggingface.co/pyannote/speaker-diarization-3.1 from PyAnnote (Bredin, 2023), which achieves state-of-the-art results (Plaquet and Bredin, 2023) across multiple datasets, to extract speaker change timestamps from the audio. In literature, the SCD is predominantly regarded as a task within the audio domain (Bredin et al., 2017), we opt not to establish an independent text-based baseline for this task. Evaluation We evaluate SCD in two ways: text-based (only valid for TokenVerse) and time-based. For both methods, predictions from TokenVerse are compared with the reference, and the F1 score is calculated. Detailed description in appendix A.
Endpointing Considering semantic endpointing, we fine-tune BERT (Devlin et al., 2019) for [ENDP] token classification on the multitask training text, termed as BERT-ENDP. Results are reported on both reference text and hypothesis text obtained from TokenVerse. From the audio perspective, we use segmentation pipeline444huggingface.co/pyannote/segmentation-3.0 from PyAnnote to obtain endpoint timestamps. Evaluation It follows the same approach as for SCD. We also report false alarms (FA), missed speech (MS), and detection error rate (DER), which are common metrics in endpointing literature (Medennikov et al., 2020).
4 Experimental Setup
@@
4.1 Datasets Descriptions
To train TokenVerse, we require conversational audio data with corresponding transcripts, NER, segment timestamps, and speaker annotations. We could not find a large-scale public dataset satisfying all the tasks. Thus, we opt for a private dataset , DefinedAI555https://www.defined.ai/. We also train and evaluate on the open-source CallHome English dataset.
@@ 4.1 Datasets Descriptions
DefinedAI contains stereo-audio/transcript pairs for contact center conversations between agents and customers. We upsampled audio from 8 kHz to 16 kHz to align with the XLSR-53 model’s requirements. Each segment includes transcripts, speaker ID and NE annotations, facilitating multitask dataset preparation. This dataset spans health, banking and finance domains, which makes it particularly challenging due to variations in NEs.
CallHome English dataset (LDC97S42) contains natural conversational stereo-audios between multiple speakers. The transcript includes named entities annotation.This dataset poses challenges due to its natural conversational nature, known to be challenging for ASR modeling, and a large number of short segments without entities, differing from the DefinedAI dataset. Further details about these datasets are provided in Table 1.
4.2 Multitask Dataset Preparation
Our work is focused on conversational audios which is typically long in duration (avg 5 minutes) and can’t be directly used for ASR training due to high GPU memory requirements. The dataset provides audio-text transcripts together with timestamp information for every segment within the long-form audio. For each sample, we begin with the first segment and find the farthest segment such that the duration is up to 20 seconds. Audios within this range are extracted as one utterance and this procedure is repeated until the last segment is consumed. Note that an utterance may span over multiple segments, potentially containing silences, noise, speaker changes, endpoints and numerous named entities. Afterward, we concatenate the text corresponding to all segments within an utterance, inserting token at appropriate positions according to our tasks, described in §2.1. This multitask dataset preparation approach applies universally across all datasets used in our experiments.
4.3 Training Details
@@
We train TokenVerse on the multitask dataset. We implement the XLSR-Transducer model from the Icefall’s Transducer recipe666https://github.com/k2-fsa/icefall/tree/master/egs/librispeech/ASR/zipformer adapted with XLSR from fairseq Ott et al. (2019). The model is optimized with pruned RNN-t loss Kuang et al. (2022). The initial learning rate is set to and we train the model for 50 epochs. For each dataset, the best epoch is selected based on WER on respective dev sets and results are presented on the eval sets. The task-transfer experiments (see §2.3) are trained for additional 10 epochs on the new task.
| Exp | Model | WER () |
| 1) | ASR (baseline) | 15.3 |
| 2) | all-tasks | 14.7 |
| 3-a) | single-[SCD] | 15.1 |
| 3-b) | single-[NE] | 14.7 |
| 3-c) | single-[ENDP] | 14.7 |
@@ 5 Results & Discussion
Automatic Speech Recognition For the DefinedAI (Tab. 2) set, including all tasks in TokenVerse (exp 2) leads to a 4% relative improvement in WER compared to the baseline ASR model (exp 1). For models trained on a single task (exp 3a-c), ASR results remain similar except for SCD. On the CallHome dataset (Tab. 5), the multitask model with all tokens yields a 7.7% relative improvement. Overall, the results on both datasets indicate that the all-tasks TokenVerse improves ASR performance.
@@
Named-Entity Recognition As expected, compared to evaluating BERT-NER on reference text, a significant degradation is observed when evaluated on hypothesis (Tab. 3) due to ASR errors Ghannay et al. (2018). In exact-match, on both the DefinedAI (Tab. 3) and CallHome (Tab. 5) test sets, the all-tasks TokenVerse outperforms the baseline BERT-NER models trained on their respective datasets and evaluated on hypothesis in F1 score. This is not the case for soft-match evaluation on the DefinedAI test set, where the F1 score is similar. This degradation is mostly attributed to the incorrect prediction of [/NE] tag by the baseline, resulting in only a partial match of the named entity words leading to increase in false positives. The absolute F1 score is low on the CallHome dataset due to higher ASR errors on named entities, attributed to their low repetition in the training text (see Tab. 1).
Speaker Change Detection On the DefinedAI (Tab. 4), including all tasks in TokenVerse outperforms the baseline PyAnnote model in time-based evaluations. Interestingly, models trained for single-task SCD perform better than the all-tasks model in terms of F1, but show similar results for Coverage-Purity based F1. Upon closer scrutiny, we found that including [ENDP] delays the prediction for [SCD] tokens, causing the hypothesis timestamps of these tokens to fall outside the tolerance window (250ms). Increasing the tolerance window further improves the F1 for both models, with a much higher rate of increase for the all-tasks model. This observation is reinforced in the text-based F1 score, where the all-tasks model achieves an F1 score of 90.3% compared to 88.5% from the single-[SCD] model. On the CallHome (Tab. 5), the all-tasks model outperforms the PyAnnote baseline. These evaluations suggest that excluding [SCD] from TokenVerse is preferable for precise speaker change timestamps, while including all tasks improves speaker-attributed text segmentation.
@@
TokenVerse Ablation Results In ASR, we observed degradation for all ablation experiments (see §2.3), with the largest relative degradation of 2.4% in WER when [ENDP] was removed. Transfer learning on any of the 3 tasks do not degrade ASR performance further. The text-based evaluations of other tasks on DefinedAI are reported in Figure 2; absolute change is calculated from the all-tasks model. Removing a task adversely affects other tasks. Specifically, for SCD and endpointing, [NE] removal has the least impact on performance. Learning it afterward either improves or maintain their performance, indicating a stronger correlation between these tasks than with NER; supported by the degradation in [SCD] performance when [ENDP] is removed. Task transfer on [ENDP] degrades the performance further, possibly due to confusion during prediction caused by the insertion of the token before [SCD] during training. Transfer to NER shows relatively large degradation compared to other tasks, likely because the model must predict both [NE] and [/NE] accurately. This suggests that tasks encoded with multiple tokens may not transfer as effectively as those encoded with a single token.
Overall, all-tasks TokenVerse outperforms specialized models for each task and single-task models suggesting that additional tasks improve each other. See sample outputs in appendix B.